OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y −2 −6 −10 −14

OpenStudy (anonymous):

@KamiBug pls help

OpenStudy (danjs):

hi

OpenStudy (danjs):

ok, so first calculate the slope, m= (y2-y) / (x2-x1)

OpenStudy (anonymous):

y= -4x-22

OpenStudy (danjs):

m = (-6-(-2)) / (2 - 1) = -4/1 = -4

OpenStudy (danjs):

y-y1 = m(x-x1) you can use any point for (x1,y1), let (x1,y1) = (1,-2) y + 2 = -4(x - 1)

OpenStudy (danjs):

y + 2 = -4x + 4 y = -4x + 2

OpenStudy (anonymous):

is that the answer and how you do it all

OpenStudy (anonymous):

thank you